# ACE-Step 1.5: Pushing the Boundaries of Open-Source Music Generation

### v1

@@ 1 Introduction
In this work, we present ACE-Step v1.5, a definitive framework designed to bridge the gap between experimental prototypes and production-ready standards. We challenge the conventional trade-off between inference efficiency and generative quality, implementing a refined Hybrid Reasoning-Diffusion Architecture that fundamentally decouples usability from acoustic rendering. Instead of an end-to-end black box, we position the Language Model (LM) as an intelligent Composer Agent and the Diffusion Transformer (DiT) as a specialized acoustic renderer. The LM functions as a high-level planner, explicitly reconstructing vague user prompts into precise structural blueprints (e.g., formatting lyrics, calculating duration, and expanding queries) to relieve the DiT of semantic ambiguities. Conversely, the DiT is freed to focus exclusively on acoustic richness and instrument separation. By integrating a novel distillation algorithm, we compress the generation process from 50 steps to merely 4–8 steps, achieving an inference speedup of over $100\times$ while simultaneously enhancing signal-to-noise performance.
@@ 5.1 Objective Metrics
Metric Sources. AudioBox scores derive from Meta’s aesthetic assessment framework; SongEval metrics are computed via the ASLP-lab evaluation suite. Style and Lyric Alignment are quantified using our proprietary ACE-Caption-Reward and ACE-Lyric-Reward models.
@@
Results. As shown in Table 1, ACE-Step 1.5 achieves the highest AudioBox CU (8.09) and PQ (8.35) scores, tied-best Coherence (4.72), and demonstrates superior temporal consistency (Mem 4.72, Nat 4.59). While commercial models like Suno-v5 maintain advantages in style/lyric alignment (46.8/34.2), ACE-Step 1.5 significantly bridges the gap compared to open-source baselines (39.1/26.3), validating that our distillation preserves high-fidelity acoustic rendering.

### v3

@@ 1 Introduction
In this work, we present ACE-Step v1.5, a definitive framework designed to bridge the gap between experimental prototypes and production-ready standards. We challenge the conventional trade-off between inference efficiency and generative quality, implementing a refined Hybrid Reasoning-Diffusion Architecture that fundamentally decouples usability from acoustic rendering. Instead of an end-to-end black box, we position the Language Model (LM) as an intelligent Composer Agent and the Diffusion Transformer (DiT [11, 8]) as a specialized acoustic renderer. The LM functions as a high-level planner, explicitly reconstructing vague user prompts into precise structural blueprints (e.g., formatting lyrics, calculating duration, and expanding queries) to relieve the DiT of semantic ambiguities. Conversely, the DiT is freed to focus exclusively on acoustic richness and instrument separation. By integrating a novel distillation algorithm, we compress the generation process from 50 steps to merely 4–8 steps, achieving an inference speedup of over $100\times$ while simultaneously enhancing signal-to-noise performance.
@@ 5.1 Objective Metrics
Metric Sources. AudioBox scores derive from Meta’s aesthetic assessment framework  [13]; SongEval metrics are computed via the ASLP-lab evaluation suite  [18]. Style and Lyric Alignment are quantified using our proprietary ACE-Caption-Reward and ACE-Lyric-Reward models.
@@
Results. As shown in Table 1, ACE-Step 1.5 achieves the highest AudioBox CU (8.09), second-best PQ (8.35), and tied-best Coherence (4.72), while demonstrating superior temporal consistency (Mem 4.72, Nat 4.59). While commercial models like Suno-v5 maintain advantages in style/lyric alignment (46.8/34.2), ACE-Step 1.5 shows strong competitiveness against open-source baselines (39.1/26.3), validating that our distillation preserves high-fidelity acoustic rendering.